OpenStudy (anonymous):

Linear algebra: I have a mysterious inner product defined for P2. ...see below...

OpenStudy (anonymous):

I've determined has the properties \[\sqrt{(x+4)* (x+4)}\] =1 \[\sqrt{(3x-5) * (3x-5)}\] =1 and (x+4) * (3x-5) = 0 Is there any way to find out with this mystery * is?